# Improved Stochastic Optimization of LogSumExp

### v2

@@ Appendix A Proofs for Section 2
By symmetry, we can see that the minimum value of the second derivative is achieved at $t^{*}=\frac{1}{2\rho}$, and it is equal to $4\rho$. Thus, for all $t\in{\rm dom}f_{\rho}$, we have that $f_{\rho}^{\prime\prime}(t)\geq 4\rho>\rho$. Thus, by (nesterov2018lectures, Theorem 2.1.11), $f_{\rho}$ is $\rho$-strongly convex. By (Zhou2018FenchelDuality, Theorem 1) this also implies that its conjugate function $f_{\rho}^{*}$ is $\frac{1}{\rho}$-smooth. ∎
@@ B.1 Related works on eOT
This subsection provides an overview of selected works on continuous entropy-regularized optimal transport. In (genevay2016stochastic), the authors tackled this problem by introducing an RKHS and optimizing the dual function (12) with SGD. This approach was extended by seguy2018large, who parameterized the dual potentials with neural networks instead of an RKHS to improve scalability. Subsequently, daniels2021score leverage this approach to approximate the optimal transport plan, using it to develop a score-based generative model. Although this direction mostly results in computationally efficient methods that works with a general cost function, a key drawback is that small values of the regularization coefficient $\varepsilon$ cause numerical instabilities due to the exponential term in the dual objective; see Remark 3.1. The work by (korotin2023neural) studies a more general formulation known as weak OT. The authors formulate it as a maximin problem and develop a neural-network-based algorithm under the assumption of a quadratic cost, a restriction that is later relaxed in (asadulaev2024neural). However, these methods are computationally intensive due to their adversarial training nature. The paper by mokrov2024energyguided approaches eOT from the perspective of energy-based models. Unfortunately, the resulting solver is computationally expensive as it involves iterative Langevin dynamics . Another popular approach to eOT in recent years is via the Schrödinger bridge (SB), e.g., (gushchin2023entropic). While SB-based solvers are also often computationally intensive, a more cost-efficient solution has been proposed by (korotin2024light). However, it relies on the quadratic cost assumption and does not support general cost. We would also like to note that a promising direction for future work is leveraging our approach for minimizing the objective (8) in (korotin2024light) to further improve scalability.
@@ B.2 Experiment with RKHS representation of dual potentials
As mentioned earlier, LSOT (seguy2018large) is inspired by the continuous eOT approach of genevay2016stochastic. This work considers a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ defined on $\mathcal{X}$, with a kernel $\kappa$, and applies SGD to solve the dual problem. Such approach suffers from the same numerical instability as LSOT, see Remark 3.1. As an alternative, we again consider the approximation (15) of the semi-dual objective which can also be maximized by SGD. Although the variable $\alpha$ is, in general, a function of $x$, we empirically found that tuning a common scalar value $\alpha\in\mathbb{R}$ for all samples works well in the experiments described below.
@@
Consider a setup analogous to the one described in Section 5 of genevay2016stochastic. Specifically, $\mu$ is a 1D Gaussian, and $\nu$ is a mixture of two Gaussians (see Figure 4 for a plot of densities). Gaussian kernel $\kappa(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{\sigma^{2}}\right)$ with a bandwidth hyperparameter $\sigma^{2}>0$ is used. The regularization coefficient is set to $\varepsilon=0.01$. We consider kernel SGD (25) applied to the dual objective as a baseline approach (genevay2016stochastic). We compare it to the proposed approach, namely, kernel SGD applied to the approximate semi-dual problem (14). For details on how the optimality gap is estimated, see Appendix B.
@@ B.3 Computing a proxy for optimality gap
    Similarly to genevay2016stochastic, we obtain a proxy $\hat{W}$ for $W(\mu,\nu)$ by solving the semi-discrete eOT problem

### v3

@@ 4 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our main goal is to make ML applications where LogSumExp minimization is needed less computationally demanding. There are many potential societal consequences of our work, yet we don’t see any direct consequences that we feel must be specifically highlighted here.
@@ Appendix A Proofs for Section 2
By symmetry, we can see that the minimum value of the second derivative is achieved at $t^{*}=\frac{1}{2\rho}$, and it is equal to $4\rho$. Thus, for all $t\in{\rm dom}f_{\rho}$, we have that $f_{\rho}^{\prime\prime}(t)\geq 4\rho>\rho$. Thus, by (Nesterov, 2018, Theorem 2.1.11), $f_{\rho}$ is $\rho$-strongly convex. By (Zhou, 2018, Theorem 1) this also implies that its conjugate function $f_{\rho}^{*}$ is $\frac{1}{\rho}$-smooth. ∎
@@ B.1 Related works on eOT
This subsection provides an overview of selected works on continuous entropy-regularized optimal transport. In (Genevay et al., 2016), the authors tackled this problem by introducing an RKHS and optimizing the dual function (13) with SGD. This approach was extended by (Seguy et al., 2018), who parameterized the dual potentials with neural networks instead of an RKHS to improve scalability. Subsequently, (Daniels et al., 2021) leverage this approach to approximate the optimal transport plan, using it to develop a score-based generative model. Although this direction mostly results in computationally efficient methods that works with a general cost function, a key drawback is that small values of the regularization coefficient $\varepsilon$ cause numerical instabilities due to the exponential term in the dual objective; see Remark 3.1. The work by (Korotin et al., 2023) studies a more general formulation known as weak OT. The authors formulate it as a maximin problem and develop a neural-network-based algorithm under the assumption of a quadratic cost, a restriction that is later relaxed in (Asadulaev et al., 2024). However, these methods are computationally intensive due to their adversarial training nature. The paper by (Mokrov et al., 2024) approaches eOT from the perspective of energy-based models. Unfortunately, the resulting solver is computationally expensive as it involves iterative Langevin dynamics . Another popular approach to eOT in recent years is via the Schrödinger bridge (SB), e.g., (Gushchin et al., 2023). While SB-based solvers are also often computationally intensive, a more cost-efficient solution has been proposed by (Korotin et al., 2024). However, it relies on the quadratic cost assumption and does not support general cost. We would also like to note that a promising direction for future work is leveraging our approach for minimizing the objective (8) in (Korotin et al., 2024) to further improve scalability.
@@ B.2 Experiment with RKHS representation of dual potentials
As mentioned earlier, LSOT (Seguy et al., 2018) is inspired by the continuous eOT approach of (Genevay et al., 2016). This work considers a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ defined on $\mathcal{X}$, with a kernel $\kappa$, and applies SGD to solve the dual problem. Such approach suffers from the same numerical instability as LSOT, see Remark 3.1. As an alternative, we again consider the approximation (16) of the semi-dual objective which can also be maximized by SGD. Although the variable $\alpha$ is, in general, a function of $x$, we empirically found that tuning a common scalar value $\alpha\in\mathbb{R}$ for all samples works well in the experiments described below.
@@
Consider a setup analogous to the one described in Section 5 of (Genevay et al., 2016). Specifically, $\mu$ is a 1D Gaussian, and $\nu$ is a mixture of two Gaussians (see Figure 4 for a plot of densities). Gaussian kernel $\kappa(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{\sigma^{2}}\right)$ with a bandwidth hyperparameter $\sigma^{2}>0$ is used. The regularization coefficient is set to $\varepsilon=0.01$. We consider kernel SGD (26) applied to the dual objective as a baseline approach (Genevay et al., 2016). We compare it to the proposed approach, namely, kernel SGD applied to the approximate semi-dual problem (15). For details on how the optimality gap is estimated, see Appendix B.
@@ B.3 Computing a proxy for optimality gap
    Similarly to (Genevay et al., 2016), we obtain a proxy $\hat{W}$ for $W(\mu,\nu)$ by solving the semi-discrete eOT problem